# Do Students Write Better Post-AI Support?

### v1

@@ Abstract
[Abstract] Academic writing increasingly involves multimodal tasks requiring students to integrate visual information and textual arguments. While generative AI (GenAI) tools, like ChatGPT, offer new pathways for supporting academic writing, little is known about how students’ GenAI literacy influences their independent multimodal writing skills or how chatbot interaction strategies (passive reactive vs. proactive scaffolding) impact learning. This study examined 79 higher education students’ multimodal academic writing performance using a comparative research design. Students completed writing tasks integrating visual data under two chatbot-assisted conditions (passive vs. proactive) and subsequently without AI assistance. Their writing performance was rigorously evaluated across five dimensions, including insightfulness, visual data integration, organisation, linguistic quality, and critical thinking. Ordinal logistic regression and correlation analyses revealed that higher levels of GenAI literacy significantly predicted stronger independent multimodal writing performance immediately after AI assistance removal, particularly for students using passive chatbots requiring active prompting. These results highlight the critical role of GenAI literacy and specific chatbot interaction strategies in shaping students’ capacities for independent multimodal academic writing. Our findings emphasise the need for purposeful integration of GenAI literacy training into curricula and balancing external scaffolding support with autonomous learning opportunities. This research offers valuable recommendations for educators leveraging AI-enhanced pedagogies to optimise student writing outcomes and technological engagement strategies.
@@
generative AI literacy, academic writing, chatbot interaction, multimodal, skill development*Practitioner Notes
What is currently known about this topic
Generative AI (GenAI) tools are increasingly used by learners in their academic writing tasks, offering new opportunities and challenges.
Effective use of advanced AI tools depends partly on students’ technological literacy and their ability to strategically engage with AI-generated content.
The academic writing process is becoming increasingly multimodal, requiring integration of textual and visual data, yet limited research exists on how GenAI tools influence learner performance in such multimodal contexts.
What this paper adds
Provides empirical evidence that learners’ GenAI literacy significantly predicts their independent multimodal academic writing performance, especially after AI supports are removed.
Demonstrates that passive chatbot interactions benefit students with higher GenAI literacy more strongly than proactive chatbot interactions, highlighting interaction style as a crucial moderating variable.
Confirms GenAI literacy as a distinct predictor of writing performance, separate from domain-specific knowledge, indicating its importance for successfully integrating visual information and critical-thinking components into AI-assisted writing tasks.
Implications for practitioners
Educators should explicitly integrate GenAI literacy training within academic curricula, focusing on skills like formulating effective prompts, critically evaluating AI-generated output, and independently integrating this assistance into academic work.
Practitioners need to carefully select chatbot scaffolding strategies appropriate for students’ GenAI literacy, balancing proactive guidance with opportunities for learners’ active inquiry, metacognitive awareness, and autonomous problem-solving.
Instructional designers should strategically balance GenAI-driven scaffolding with independent practice opportunities for students to independently apply and develop their GenAI literacy skills.
1 Introduction
@@
Academic writing is an essential practice in higher education. It helps students develop critical thinking, effective communication, and the ability to deal with complex ideas and information 33, 46. Students need to interpret and synthesise data, create logical arguments, and assess different perspectives 13, 47. Increasingly, academic writing also involves multimodal literacy, requiring students to accurately interpret, synthesise, and incorporate varied visual data sources, such as graphs, charts, or images, into their arguments. This skill reflects the increasingly multimodal nature of contemporary academic discourse and demands additional capabilities concerning visual-data integration 34, 46, 27. Academic writing is crucial for both academic success and professional competency 53, 59. Employers value strong academic writing skills as they help individuals write persuasive reports, create coherent presentations, and solve problems effectively 12. Yet, academic writing is linked to cognitive and metacognitive processes. It encourages students to monitor and control their thought processes, facilitating metacognitive awareness and self-regulation 18. This engagement is essential for students’ ability to plan, evaluate, and reflect on their learning strategies, leading to a deeper understanding and better retention of information 21.
@@
As technology advances, academic writing can evolve from being solely an individual process to one supported by various tools 45. With the increasing availability of generative AI (GenAI) tools (e.g., ChatGPT and Gemini), students have more opportunities to enhance their academic writing performance 62, 4. Using GenAI tools in writing is becoming more common 30, making it urgent to understand how these external supports impact students’ skill development 39. GenAI tools can provide feedback, generate content ideas, or assist in structuring arguments, offering new ways for students to engage with writing tasks 7. However, while these tools present exciting possibilities, it is critical to examine their influence on learners’ development of independent writing skills 63. Understanding this dynamic will guide educators in integrating technology effectively into writing instruction, ensuring that while students benefit from technological aids, they also continue to grow their critical thinking and independent writing abilities 45, 7.
Indeed, emerging studies are beginning to explore how GenAI can produce high-quality writing and support student writing processes 38, 30, 39. Most of these studies focus on how GenAI can assist humans in writing and the patterns that emerge from this interaction 45, 62. However, existing research predominantly addresses text-only writing contexts, leaving uncertainties about the role of GenAI in supporting multimodal academic writing involving visual information 27, 34. There is also little empirical evidence on how GenAI tools affect students’ academic writing performances, especially when these tools are taken away after they have been used. Using GenAI tools might merely create the illusion of performance enhancement, but it might not necessarily lead to learning and skill improvements 16, 63. Additionally, few studies have thoroughly examined how different interaction strategies of GenAI, such as passively replying versus proactively scaffolding, impact student writing. Understanding interaction strategy is important because it could shape how students engage with these tools and influence the development of their writing skills 45, 29. A passive interaction might mean the tool only responds to commands, while proactively scaffolding could guide students step-by-step and provide feedback 63. Each style can affect how independent and critical students become in their writing. Thus, it is crucial to understand if dependence on GenAI affects students’ writing skills and independence. Research in this area can help educators balance the use of GenAI technology, making sure students can take advantage of GenAI while still growing their own academic writing abilities 45, 16, 7.
The current study contributes empirical evidence to address the aforementioned gaps. Specifically, we conducted a comparative study with 79 higher education students who completed an academic writing task based on visual data, with the support of a passive (reactive replying) or proactive (active scaffolding) GenAI agent. Additionally, students completed a similar multimodal academic-writing task with visual information integration but with GenAI support removed, to examine potential lasting effects. Students’ written responses were rigorously scored using a comprehensive rubric covering insightfulness, visual data integration, organisation and coherence, linguistic quality, and critical thinking. Furthermore, we investigated how students’ generative AI literacy influenced their multimodal academic writing performance when provided with and subsequent to the removal of GenAI assistance. By addressing these points, our study contributes novel insights into the differential impacts of passive versus proactive interaction strategies on students’ abilities to integrate textual and visual data, and underscores the role of learners’ varying GenAI literacy in determining academic writing performances. These findings extend existing research by offering practical implications for educators and practitioners aiming to integrate GenAI effectively into multimodal academic writing practices.
@@ 2 Background
2.1 Academic Writing
Academic writing is a core skill in higher education, requiring students to effectively synthesise ideas, engage critically with evidence, and communicate coherently. At the heart of successful academic writing lies the ability to provide insightful interpretations by deeply analysing the subject matter, demonstrating originality, and making sophisticated connections among multiple sources 32, 59. Beyond text alone, contemporary academic discourse increasingly involves multimodal literacy, in particular through the integration of visual data. Students must accurately interpret, meaningfully synthesise, and seamlessly incorporate varied graphical sources into their narrative 34, 46, 27. Equally central is the structural organisation and coherence of written arguments, allowing ideas to flow logically and enhancing reader comprehension 23, 53. Further, linguistic quality, marked by appropriate grammar, syntax, vocabulary, and academic style, is indispensable, ensuring precision, clarity and credibility within scholarly communication 5, 24. Lastly, academic writing requires rigorous critical thinking, encompassing analytical evaluation, reasoned argumentation, and thorough exploration of competing ideas or solutions 43, 15.
Assessing students’ academic writing holistically thus demands evaluation across these multiple dimensions. Insightfulness is typically measured through evidence of original, nuanced interpretations rather than superficial summaries 32, 9. Visual data integration is judged by examining the accuracy and relevance of references to graphical material, as well as how effectively visual information enhances overall arguments 27. Organisation and coherence involve assessing logical structuring and the clarity with which ideas connect and flow progressively 47. Linguistic quality is commonly evaluated through accuracy in grammar, usage, and stylistic appropriateness, reflecting scholarly conventions and readability 5. Finally, students’ critical thinking skills are commonly demonstrated through their ability to carefully analyse evidence, balance perspectives, and propose reasoned solutions or arguments based on comprehensive evaluation 15, 43. Combined, these dimensions offer a nuanced, comprehensive assessment of students’ academic writing performance, which is important for capturing students’ academic success and intellectual growth.
2.2 Generative AI and Writing
[…]
@@ 2.3 Generative AI Literacy
Given the increasing use of GenAI in educational contexts, particularly in academic writing, it has become essential to investigate and consider students’ GenAI literacy when exploring AI-assisted writing practices. While general AI literacy has been defined and explored extensively in previous frameworks 40, 44, the unique characteristics of generative technologies necessitate a more specialised literacy framework. Researchers have highlighted that GenAI literacy involves not only understanding how generative tools function, but also mastering practical skills to effectively prompt these tools, critically evaluate AI-generated content, and navigate ethical and contextual complexities specific to GenAI-generated outputs 6, 69. Despite recognition of the importance of GenAI literacy for meaningful interaction with tools like ChatGPT and Gemini, there remains limited empirical research that explicitly examines how students’ varying levels of GenAI literacy impact their academic writing performances when GenAI assistance is provided or withdrawn 3, 41. Addressing this gap is imperative, as understanding the relationship between students’ GenAI literacy and their academic writing outcomes can inform practical strategies for educators seeking to effectively integrate these advanced technologies into multimodal academic writing instruction 11, 3.
@@ 2.4 Domain Knowledge
Domain knowledge is an essential consideration when investigating AI-assisted multimodal academic writing, particularly when visual information and interpretation are integral to the task. Students’ capacity to successfully integrate visual data into meaningful written discussions depends heavily on their familiarity with relevant subject areas 1, 55. This underscores the necessity of evaluating and accounting for the domain knowledge possessed by individual learners, especially when examining the relationship between their GenAI literacy and writing performance 2. Without considering domain proficiency, researchers may inadvertently attribute students’ challenges in producing coherent multimodal arguments or effectively utilising GenAI supports to deficiencies in GenAI literacy alone, rather than identifying underlying gaps in subject-specific knowledge 1. Accordingly, research into GenAI literacy for multimodal academic writing contexts should adopt a nuanced approach, explicitly acknowledging the role of domain-specific expertise alongside GenAI literacy.
@@ 2.5 Research Gaps
The current study aimed to address the aforementioned gaps by examining how students’ GenAI literacy affects their multimodal academic writing performance with and without GenAI support. Specifically, we investigated whether these effects differ based on chatbot interaction strategies (passive vs. proactive) and account for students’ domain-specific knowledge to ensure an accurate interpretation of the effects of GenAI literacy. Specifically, the following research questions were investigated:
RQ1: What is the relationship between students’ GenAI literacy and their academic writing performance at the AI-support and AI-removal stages, and does this relationship differ based on chatbot interaction type (passive vs. proactive)?
RQ2: Do students with low versus high GenAI literacy differ significantly in their academic writing performance at the AI-support and AI-removal stages, and do these differences differ based on chatbot interaction type (passive vs. proactive)?
RQ3: To what extent does students’ GenAI literacy predict their academic writing performance at the AI-removal stage after controlling for domain-specific knowledge, chatbot interaction type (passive vs. proactive), and writing performance at the AI-support stage?
@@ 3.1 Learning Contexts
The learning context involved writing based on visual information, a critical skill in academic writing, closely aligned with tasks such as IELTS Academic Writing Task 1 25. Participants received three distinct visual representations (a bar chart, a network diagram, and a ward map; described in detail in Section 3.3.1). They were instructed to interpret these visuals and compose a short written answer (150 words) addressing a question about team performance (e.g., "How the two nurses managed the primary patient (Amy) while attending to other beds, focusing on their task prioritisation, verbal communication, and stress levels?"). The visuals were specifically designed using data from authentic learning contexts, such as healthcare simulations, ensuring participants had no prior exposure to these materials. Thus, any potential effects stemming from participants’ prior familiarity with the presented information were minimal. The study was conducted through an online environment (details in Section 3.3.2).

### v2

# The Agency Gap: How Generative AI Literacy Shapes Independent Writing after AI Support

@@ Abstract
Generative AI (GenAI) tools are rapidly transforming higher education, yet little is known about how students’ GenAI literacy shapes their ability to perform independently once such support is removed. This study investigates what we term the agency gap, introduced as the extent to which GenAI literacy predicts student writing performance in contexts that require self-initiation and regulation. Seventy-nine medical and nursing students completed multimodal academic writing tasks based on visual data, supported either by a reactive or proactive GenAI chatbot, followed by a parallel task without AI support. Writing was evaluated across insightfulness, visual data integration, organisation, linguistic quality, and critical thinking. Results showed that GenAI literacy predicted independent writing performance only in the reactive condition, where students had to actively mobilise their own strategies. Mediation analyses revealed no indirect effect via in-task performance, indicating that GenAI literacy acts as a direct, task-general competence rather than a proxy for domain knowledge or other literacies. By contrast, proactive scaffolding equalised outcomes across literacy levels, reducing reliance on learners’ GenAI literacy. The agency gap highlights when GenAI literacy matters most, with implications for designing equitable AI-supported learning environments that both leverage and mitigate differences in students’ GenAI literacy.
@@
[…]
@@
Academic writing is an essential practice in higher education. It helps students develop critical thinking, effective communication, and the ability to deal with complex ideas and information (Lea and Street, 1998; Oshima and Hogue, 2007). Students need to interpret and synthesise data, create logical arguments, and assess different perspectives (Durst, 1987; Paltridge, 2004). Increasingly, academic writing also involves multimodal literacy, requiring students to accurately interpret, synthesise, and incorporate varied visual data sources, such as graphs, charts, or images, into their arguments. This skill reflects the increasingly multimodal nature of contemporary academic discourse and demands additional capabilities concerning visual-data integration (Lea and Street, 2006; Oshima and Hogue, 2007; Jewitt, 2006). Academic writing is crucial for both academic success and professional competency (Swales and Feak, 1994; Wingate, 2012). Employers value strong academic writing skills as they help individuals write persuasive reports, create coherent presentations, and solve problems effectively (Council, 2012). Yet, academic writing is linked to cognitive and metacognitive processes. It encourages students to monitor and control their thought processes, facilitating metacognitive awareness and self-regulation (Flavell, 1979). This engagement is essential for students’ ability to plan, evaluate, and reflect on their learning strategies, leading to a deeper understanding and better retention of information (Hacker et al., 2009).
@@
As technology advances, academic writing can evolve from being solely an individual process to one supported by various tools (Nguyen et al., 2024). With the increasing availability of generative AI (GenAI) tools (e.g., ChatGPT and Gemini), students have more opportunities to enhance their academic writing performance (Yan, 2023; Bašić et al., 2023). Using GenAI tools in writing is becoming more common (Kim et al., 2025), making it urgent to understand how these external supports impact students’ skill development (Liu et al., 2024). GenAI tools can provide feedback, generate content ideas, or assist in structuring arguments, offering new ways for students to engage with writing tasks (Cardon et al., 2023). Yet their growing use raises a critical question: do GenAI tools genuinely build students’ independent writing skills, or do they merely provide temporary performance boosts (Fan et al., 2025; Yan et al., 2024a)? This issue is especially salient in multimodal contexts, where students are required to write while interpret and synthesise visual information (Liu et al., 2024). While research has begun to explore how GenAI assists writing (Nguyen et al., 2024), little is known about whether its benefits transfer once the technology is removed, or how different interaction strategies, such as reactive replying versus proactive scaffolding, shape that transfer (Jin et al., 2025).
A further underexplored dimension is the role of students’ GenAI literacy. Beyond general academic or domain-specific literacies, GenAI literacy encompasses knowledge of how AI systems operate, the ability to formulate effective prompts, the capacity to critically evaluate outputs, and an awareness of associated risks and ethical considerations (Jin et al., 2024; Bozkurt, 2024; Ng et al., 2021). Students with higher levels of GenAI literacy may use GenAI tools more effectively while also sustain their performance once such support is removed. Yet it remains unclear whether GenAI literacy acts indirectly, by improving AI-assisted work that then transfers, or directly, as a transferable competence that enables independent performance.
In this study, we investigate what we term the agency gap—the extent to which GenAI literacy predicts independent academic writing when learners must self-initiate and regulate their engagement with AI. Seventy-nine higher education students completed multimodal academic writing tasks with either reactive (responding only to queries) or proactive (actively scaffolding) GenAI support, followed by a parallel task without AI. Results showed that GenAI literacy predicted independent writing performance only in the reactive condition, where learner agency was essential, and not in the proactive condition, where scaffolding equalised outcomes across literacy levels. Mediation analyses further revealed no indirect effect via AI-supported performance, indicating that GenAI literacy acts as a direct, task-general competence rather than a proxy for domain knowledge or writing proficiency. By conceptualising and evidencing the agency gap, this study advances understanding of when and how GenAI literacy matters and highlights two complementary levers for practice: cultivating GenAI literacy to support independent performance, and designing proactive scaffolds to mitigate disparities when learner agency cannot be assumed.
@@ 2 Background
2.1 GenAI and Academic Writing
Academic writing is a core skill in higher education, requiring students to synthesise ideas effectively, engage critically with evidence, and communicate coherently. At its heart, strong writing demonstrates insightfulness, where students provide original and nuanced interpretations by deeply analysing subject matter and connecting ideas across sources (Kuhn, 2005; Wingate, 2012). Contemporary academic discourse also increasingly requires visual data integration, as students are increasingly expected to interpret and incorporate charts, diagrams, and other graphical information into their arguments (Lea and Street, 2006; Oshima and Hogue, 2007; Jewitt, 2006). Equally essential is organisation and coherence, ensuring that arguments flow logically and ideas are structured in ways that support reader comprehension (Hyland, 2021; Swales and Feak, 1994). The quality of academic writing further relies on linguistic quality, reflected in precise grammar, appropriate vocabulary, and adherence to academic style conventions (Biber and Gray, 2010; Hyland and Milton, 1997). Finally, rigorous critical thinking underpins effective writing by enabling students to evaluate evidence, consider alternative perspectives, and develop reasoned arguments (Moon, 2007; Facione, 2000). A holistic assessment of academic writing, therefore, demands comprehensive evaluation across all of the above dimensions, to capture the multifaceted nature and reflect the complexity student face in authentic academic writing.
More recently, the advancements in GenAI have transformed writing tools into sophisticated assistants for diverse writing tasks (Lee et al., 2024). Research has produced systems for AI-powered writing support, such as CoAuthor, which provides sentence-level suggestions integrated into the writing process (Lee et al., 2022), and tools that generate text continuations based on user input (Jakesch et al., 2023). Studies show that students expect GenAI to serve multiple roles, including as an assistant, tutor, and collaborator (Kim et al., 2025), though they also have concerns about its ethical and practical implications (Wang, 2025). Further work has explored the behavioural patterns that emerge during GenAI-assisted writing (Shibani et al., 2023; Nguyen et al., 2024), finding that iterative and interactive engagement correlates with improved outcomes (Nguyen et al., 2024) and that engagement in higher-order cognitive processes is linked to greater textual sophistication (Yang et al., 2025). Despite these advances, current research primarily addresses text-based writing, leaving GenAI’s role in multimodal writing that involves visual data integration largely unexplored. Furthermore, the relationships between students’ GenAI literacy and their subsequent writing performance remain under-investigated, limiting the ability to offer tailored, literacy-specific support. Finally, little research has examined how different chatbot interaction strategies, such as reactively responding versus proactively scaffolding, impact academic writing. This is a critical gap, as interaction strategies significantly shape student engagement and learning outcomes during human-AI collaboration (Nguyen et al., 2024; Jin et al., 2025).
2.2 GenAI Chatbot Interaction Strategies
The integration of GenAI chatbots, particularly conversational chatbot, into educational technologies has shown a growing trend as a solution to offer scalable and dynamic support to empower existing educational technologies (Akpan et al., 2024; Bozkurt, 2023). Conversational chatbots are designed to infuse interactivity into learning environments, providing explanations, clarifications and insights that enhance students’ cognitive and emotional engagement (Wu et al., 2025). The most widely adopted form of GenAI chatbot in education is the conventional reactive chatbot. These chatbots behave in a reactive manner, responding to student-initiated queries by delivering contextually relevant information. Their value lies in providing clarity and explanation when prompted, which can effectively support student understanding in knowledge-intensive tasks. Empirical studies have demonstrated the use of reactive chatbots can enhance learning across varied contexts and populations. They have been found to support knowledge-based dialogue and boost creative problem-solving performance among postgraduate students (Song et al., 2025), increase motivation and reduce stress while enabling self-paced learning in secondary programming education (Ait Baha et al., 2023), and improve science knowledge, engagement, and motivation when compared with rule-based systems (Ng et al., 2024). However, despite these benefits, reactive chatbots require students equipped with ability to initiate effective queries, meaning that learners with limited GenAI literacy may struggle to maximise their value.
Recent work has moved beyond the limitations of reactive chatbots toward the design of proactive scaffolding chatbots informed by educational theory. Scaffolding, rooted in Vygotsky’s zone of proximal development, supports learners by breaking complex tasks into smaller steps, asking guiding questions, and providing tailored feedback that promotes self-regulation (Gibbons, 2002). By operationalising these principles, proactive chatbots engage learners through structured narratives, dynamically adapt to student input, and reduce reliance on prompting skills, thereby providing more equitable support. Empirical studies have demonstrated their effectiveness across diverse contexts. For instance, they have improved comprehension of complex visualisations by guiding step-by-step interpretation (Yan et al., 2025), enhanced instructional design performance through structured exploration and feedback (Bai et al., 2024), facilitate personalised learning through structured tutor–student dialogues that enhance knowledge acquisition and critical analysis (Looi and Jia, 2025), and support self-regulated strategy development in reading through interactive guidance that enhanced metacognitive and motivational engagement (Pan et al., 2025). However, there is still limited research on how different chatbot interaction strategies shape students’ writing performance, particularly in relation to differences in their levels of GenAI literacy.
@@ 2.3 Generative AI Literacy
Given the increasing use of GenAI in educational contexts, particularly in academic writing, it has become essential to investigate and consider students’ GenAI literacy when exploring AI-assisted writing practices. While general AI literacy has been defined and explored extensively in previous frameworks (Long and Magerko, 2020; Ng et al., 2021), the unique characteristics of generative technologies necessitate a more specialised literacy framework. Researchers have highlighted that GenAI literacy involves not only understanding how generative tools function, but also mastering practical skills to effectively prompt these tools, critically evaluate AI-generated content, and navigate ethical and contextual complexities specific to GenAI-generated outputs (Zhao et al., 2024; Bozkurt, 2024). Despite recognition of the importance of GenAI literacy for meaningful interaction with tools like ChatGPT and Gemini, there remains limited empirical research that explicitly examines how students’ varying levels of GenAI literacy impact their academic writing performances when GenAI assistance is provided or withdrawn (Annapureddy et al., 2024; Lyu et al., 2024). Addressing this gap is imperative, as understanding the relationship between students’ GenAI literacy and their academic writing outcomes can inform practical strategies for educators seeking to effectively integrate these advanced technologies into multimodal academic writing instruction (Chiu, 2024; Annapureddy et al., 2024).
@@ 2.4 Domain Knowledge
Domain knowledge is an essential consideration when investigating AI-assisted multimodal academic writing, particularly when visual information and interpretation are integral to the task. Students’ capacity to successfully integrate visual data into meaningful written discussions depends heavily on their familiarity with relevant subject areas (Alexander, 1992; Tricot and Sweller, 2014). This underscores the necessity of evaluating and accounting for the domain knowledge possessed by individual learners, especially when examining the relationship between their GenAI literacy and writing performance (Alexander and Judy, 1988). Without considering domain proficiency, researchers may inadvertently attribute students’ challenges in producing coherent multimodal arguments or effectively utilising GenAI supports to deficiencies in GenAI literacy alone, rather than identifying underlying gaps in subject-specific knowledge (Alexander, 1992). Accordingly, research into GenAI literacy for multimodal academic writing contexts should adopt a nuanced approach, explicitly acknowledging the role of domain-specific expertise alongside GenAI literacy.
@@ 2.5 Research Gaps
The current study aimed to address the aforementioned gaps by examining how students’ GenAI literacy affects their academic writing performance with and without GenAI support. Specifically, we investigated whether these effects differ based on chatbot interaction strategies (reactive vs. proactive) and account for students’ domain-specific knowledge to ensure an accurate interpretation of the effects of GenAI literacy. The following research questions were investigated:
RQ1: To what extent does students’ GenAI literacy relate to their academic writing performance, after controlling for their visualisation literacy, and how does this relationship differ across chatbot interaction types (reactive vs. proactive) and at the AI-support versus AI-removal stages?
RQ2: Do students with low versus high GenAI literacy differ significantly in their academic writing performance at the AI-support and AI-removal stages, and do these differences differ based on chatbot interaction type (reactive vs. proactive)?
RQ3: To what extent does students’ GenAI literacy predict their academic writing performance at the AI-removal stage after controlling for domain-specific knowledge, chatbot interaction type (reactive vs. proactive), and writing performance at the AI-support stage?
RQ4: To what extent does students’ academic writing performance during the AI-support stage mediate the relationship between their GenAI literacy and their performance at the AI-removal stage?
@@ 3.1 Learning Contexts
The learning context involved writing based on visual information, a critical skill in academic writing, closely aligned with tasks such as IELTS Academic Writing Task 1 (IELTS, 2025). Participants received three distinct visual representations (a bar chart, a network diagram, and a ward map; described in detail in Section 3.3.1). They were instructed to interpret these visuals and compose a short written answer (150 words) addressing a question about team performance (e.g., "How the two nurses managed the primary patient (Amy) while attending to other beds, focusing on their task prioritisation, verbal communication, and stress levels?"). The visuals were specifically designed using data from authentic learning contexts, such as healthcare simulations, ensuring participants had no prior exposure to these materials. Thus, any potential effects stemming from participants’ prior familiarity with the presented information were minimal. The study was conducted through an online environment (details in Section 3.3.2).